# Mitigating Metropolitan Carbon Emissions with Dynamic Eco-driving at Scale

### v1

@@ 6.3 Impact Assessment
Impact assessment involves leveraging the trained eco-driving policies to assess the benefits of eco-driving in the representative scenario set $\Phi^{\prime}$. Let $\Pi$ be the trained policy set for all 28 partitions. We further employ the zero-shot generalizability of these policies to conduct the final impact assessment. In particular, for a given semi-autonomous vehicle penetration level, we have seven partitions by varying the lane count and phase count. Recall that we denote each partition as $\psi_{i}$ where $i\in\{1,\cdots,7\}$ when the semi-autonomous vehicle penetration level is given. Let $\Pi^{{}^{\prime}}=[(\pi_{\psi_{i}}^{1},\pi_{\psi_{i}}^{2})_{i\in\{1,\cdots,7\}}\cup\pi^{b}$] be the subset of trained policies for all seven partitions under a given semi-autonomous vehicle penetration level and the status quo driving baseline control policy $\pi^{b}$.
For a given scenario $\phi$ belonging to one of the seven clusters, we seek the best-performing policy $\pi^{*}_{\phi}$ as follows,
$\pi^{*}_{\phi}=\operatorname*{argmin}_{\pi^{k}\in\Pi^{{}^{\prime}}}(g(\pi^{k},\pi^{b},\phi))$ (15)
@@
$g(\pi^{k},\pi^{b},\phi)=\begin{cases}f(\pi^{k},\phi)&\text{if network decomposition constraints are satisfied}\\ M&\text{otherwise}\end{cases}$ (16)
such that $f(\pi,\phi)$ denoting the function that captures the CO2 emission for the scenario $\phi$ under the control law $\pi$ and $M$ is large enough constant to indicate the policy $\pi^{k}$ is invalid due to any of the network decomposition constraint violations as listed in Equations 6, 7, and 8. If any of the constraints are not satisfied, we default to the status quo driving baseline control policy $\pi^{b}$. Below, we describe how we check for these constraint violations.
Control active intersection throughput constraint: To check if the control policy $\pi^{k}$ satisfies the throughput requirement, we compare its throughput with the status quo driving baseline $\pi^{b}$ throughput. If $n_{\pi^{k}}$ is the throughput of $\pi^{k}$ and $n_{\pi^{b}}$ is their throughput of $\pi^{b}$, then the constraint is satisfied if $n_{\pi^{k}}\geq n_{\pi^{b}}$.
@@
Preceding intersection vehicle waiting time constraint: When eco-driving is performed at the control active intersection, the changed vehicle flow may affect the outflow of the intersection immediately precedes it (e.g., when the vehicles glide at the control active intersection slowing the rate other vehicles leave the preceding intersection). To ensure this does not happen, we check if $w_{\pi^{k}}\leq w_{\pi^{b}}$ where $w_{\pi^{k}}$ and $w_{\pi^{b}}$ are the average waiting time per vehicle in the preceding intersection under eco-driving policy $\pi^{k}$ and status quo driving baseline $\pi^{b}$, respectively. In simulations, this also includes any vehicle that has been scheduled to be spawned in the road network but yet to be spawned due to congested traffic settings.
Finally, by performing above zero-shot transfer scheme to all scenarios in $\Phi^{{}^{\prime}}$, we obtain the best policy $\pi^{*}_{\phi}$ for each scenario $\phi$. We then use the best policy of each scenario for subsequent impact assessments. In particular, regional eco-driving effectiveness $E(\cdot)$ is used to assess the overall benefit level as per Equation 4. Each scenario is simulated five times for statistical significance, and results are averaged following the recommended practice Agarwal et al. (2021).
@@ 7.1 Experimental setup
We use Proximal Policy Optimization (PPO) Schulman et al. (2017) algorithm in training our MT-DRL policies with actor-critic architecture. The actor and critic do not share parameters and are separate neural networks. Centralized training and a decentralized execution paradigm are used for training. Section 10.8 in the Appendix describes the architecture details and all the hyperparameter configurations used in training.
@@ 7.2 Impact of eco-driving
Our eco-driving assessment encompasses diverse perspectives. First, we analyze the overall regional eco-driving effectiveness and the resultant variations in the intersection throughput in the three cities. Second, we analyze which intersection-related factors contribute greater influence at different levels of eco-driving adoption. Third, we analyze the level of intersection compatibility needed to achieve the benefits, informing implementation plans. Fourth, we analyze eco-driving behaviors and examine the safety implications of eco-driving, an aspect that has not received adequate attention in previous work Carslaw et al. (2010). Last, we contextualize eco-driving by analyzing its role in joint consideration with travel growth projections and decarbonization technologies such as vehicle electrification and the adoption of hybrid vehicles.
7.2.1 Eco-driving effectiveness
Figures 8 show the annual average effectiveness of eco-driving, assessing the percentage emission reduction and average throughput improvement. Eco-driving achieves a notable 11-22% emission reduction, equivalent to the national carbon emissions of Israel and Nigeria, respectively (refer to Appendix 10.2 for the estimation). Notably, a 10% eco-driving adoption can yield 25-50% of the total emission reduction, showcasing a non-linear benefit scaling and greater potential for short-term deployments to gain a majority of total benefits. Importantly, these gains incur no costs to intersection throughput. Additionally, Figure 9 demonstrates consistent emission benefits under different weather conditions across seasons, with minor fluctuations. This emphasizes that eco-driving can yield benefits throughout the year.
@@ 7.2.1 Eco-driving effectiveness
Figure 8: Left Annual average emission reduction vs. eco-driving adoption (by factoring all seasons and peak and off-peak hours). Right Throughput improvement with varying adoption levels. By design, the throughput improvement is at least 0% when evaluating the emission reduction. Also, we ensure there throughput increases are not negatively affect the network level traffic.
Figure 9: Impact of weather conditions on eco-driving benefits. For each city, under every season, we present the eco-driving emission benefits with two different weather conditions: sunny and rainy. Not all cities go through every weather condition for a given season. We shift some of the curves within ±$\pm 2$% on the x-axis for better visibility.
@@ 7.2.2 Influential factors in eco-driving
Through a Pearson correlation analysis in Figure 11, we underscore key factors influencing emission benefits for informed intersection design and city planning. We consider nine key eco-driving factors that are known to have a significant impact on emission benefits and are directly observable, including temperature, humidity, speed limit, road grade, signal time ratio, vehicle inflow, phase count, lane count, and lane length. The signal ratio denotes the ratio between approach-related phase times and total cycle time. Notably, speed limits, vehicle inflow, and lane lengths significantly impact emission benefits. In particular, when the speed limit and lane length increase, the benefit levels increase. On the other hand, when the inflow increases, the benefit decreases. This explains why the three cities we consider have different benefit levels. As discussed in Appendix Section 10.5, different cities have different intersection feature distributions (refer to Figure 17 in the Appendix). Thus, cities like Atlanta demonstrate higher benefits, given their higher speed limits and longer incoming approaches, compared to relatively low-benefit cities like San Francisco.
@@
7.2.3 Intersection compatibility
Figure 12 shows Pareto charts for all three cities revealing that at least 70% of emission benefits can be achieved by implementing eco-driving in just 20% of intersections at each adoption level. This emphasizes the potential for an efficient phased deployment, focusing on a specific subset of intersections. However, the Venn diagram in Figure 13 illustrates the specific 20% of intersections delivering the most significant emission benefits at every adoption level may have little overlap between eco-driving adoption levels. It highlights that intersections effective at lower eco-driving adoption levels may not maintain the same efficacy as eco-driving adoption increases, emphasizing the need for careful deployment planning, such as which intersections to be embedded with roadside units (RSU) for Signal and Phase Timing (SPaT) communication. In broad terms, this would question what infrastructure is considered long-term value for eco-driving as transportation agencies are often constrained on monetary budgets for infrastructure.
@@ 7.2.4 Eco-driving behaviors
Figure 14: Time-space diagrams of semi-autonomous vehicles (highlighted) and human-driven vehicles (muted) for a selected representative two-lane intersection.
We further analyze the induced road safety metrics of these learned behaviors compared to human-like driving. Figure 15 presents a roadway safety analysis using widely used surrogate safety measures (SSMs) commonly employed in assessing roadway safety Wang et al. (2021a). We analyze five SSMs, namely: time to collision (TTC), post-encroachment time (PET), maximum speed (MaxS), change in speed (DeltaS), and deceleration rate (DR). Further details on the analysis and used surrogate measures are in Appendix Section 10.9. Results in Figure 15 suggest that the learned behaviors are as safe as human-like driving, resulting from smoother vehicle behaviors. However, given that eco-driving behaviors of controlled vehicles are emerging behaviors that may come as counterintuitive for regular human drivers, it may trigger unforeseen human-driver responses, requiring additional human-factor studies for a better understanding of such human factors at play Jayawardana et al. (2023).
@@
Figure 16: Eco-driving benefits in the presence of vehicle electrification and hybrid engines as alternative decarbonization strategies. Figure shows total reductions over 2024-2050 in carbon emissions at various constant eco-driving adoption levels amidst projected electric vehicle and hybrid adoption in each city. Eco-driving yields additional benefits relative to a future with hybridization or electrification alone. For example, with an average total adoption rate of 20% from 2024 to 2050, eco-driving alone would cut San Francisco emissions in that period by about 7%, and when combined with projected eclectic vehicle and hybrid adoption, would cut emissions by about 17%. Results are relative to an all internal combustion engine vehicles baseline (i.e., without electrification or hybrid adoption.)
@@ 9 Acknowledgments
This research was partially supported by the Utah Department of Transportation (Project F-ST99(783)), and the authors extend their appreciation to Michael Sheffield, Christopher Siavrakas, and Kelly Njord at the Utah Department of Transportation for constructive discussions and providing their expert opinions on the work. Authors also extend their appreciation to Steven Barrett, Samitha Samaranayake, Zuduo Zheng, Yashar Farid, Jinhua Zhao, and Felice Frankel for their constructive comments. The authors also thank Sunera Avinash, Anirudh Valiveru, Jiaxin He, and Dajiang Suo for their help in numerous stages of the project. The authors acknowledge MIT SuperCloud and the Lincoln Laboratory Supercomputing Center for providing computational resources supporting the research results in this paper.

### v2

@@ 6.3 Impact Assessment
Impact assessment involves leveraging the trained eco-driving policies to assess the benefits of eco-driving in the representative scenario set $\Phi^{\prime}$. Let $\Pi$ be the trained policy set for all 28 partitions. We further employ the zero-shot generalizability of these policies to conduct the final impact assessment. In particular, for a given semi-autonomous vehicle penetration level, we have seven partitions by varying the lane count and phase count. Recall that we denote each partition as $\psi_{i}$ where $i\in\{1,\cdots,7\}$ when the semi-autonomous vehicle penetration level is given. Let $\Pi^{{}^{\prime}}=[(\pi_{\psi_{i}}^{1},\pi_{\psi_{i}}^{2})_{i\in\{1,\cdots,7\}}\cup\pi^{b}$] be the subset of trained policies for all seven partitions under a given semi-autonomous vehicle penetration level and the status quo driving baseline control policies $\pi^{b}\coloneqq(\pi^{b}_{\phi})_{\phi\in\Phi^{\prime}}$.
For a given scenario $\phi$ belonging to one of the seven clusters, we seek the best policy $\pi^{*}_{\phi}$,
$\pi^{*}_{\phi}=\operatorname*{argmin}_{\pi^{k}\in\Pi^{{}^{\prime}}}(g(\pi^{k},\pi^{b}_{\phi},\phi))$ (17)
@@
$g(\pi^{k},\pi^{b}_{\phi},\phi)=\begin{cases}f(\pi^{k},\phi)&\text{if network decomposition constraints are satisfied}\\ M&\text{otherwise}\end{cases}$ (18)
such that $f(\pi^{k},\phi)$ denoting the function that captures the CO2 emission for the scenario $\phi$ under the control law $\pi^{k}$ and $M$ is large enough constant to indicate the policy $\pi^{k}$ is invalid due to any of the network decomposition constraint violations as listed in Equations 8, 9, and 10. If any of the constraints are not satisfied, we default to the status quo driving baseline control policy $\pi^{b}$. Below, we describe how we check for these constraint violations.
Control active intersection throughput constraint: To check if the control policy $\pi^{k}$ satisfies the throughput requirement, we compare its throughput with the status quo driving baseline $\pi^{b}$ throughput. If $n_{\pi^{k}}$ is the throughput of $\pi^{k}$ and $n_{\pi^{b}}$ is the throughput of $\pi^{b}$, then the constraint is satisfied if $n_{\pi^{k}}\geq n_{\pi^{b}}$.
@@
Preceding intersection vehicle waiting time constraint: When eco-driving is performed at the control active intersection, the changed vehicle flow may affect the outflow of the intersection immediately precedes it (e.g., when the vehicles glide at the control active intersection, slowing the rate other vehicles leave the preceding intersection). To ensure this does not happen, we check if $w_{\pi^{k}}\leq w_{\pi^{b}}$ where $w_{\pi^{k}}$ and $w_{\pi^{b}}$ are the average waiting time per vehicle in the preceding intersection under eco-driving policy $\pi^{k}$ and status quo driving baseline $\pi^{b}$, respectively. In simulations, this also includes any vehicle that has been scheduled to be spawned in the road network but yet to be spawned due to congested traffic settings.
Finally, by performing the above zero-shot transfer scheme to all scenarios in $\Phi^{{}^{\prime}}$, we obtain the best policy $\pi^{*}_{\phi}$ for each scenario $\phi$. We then use the best policy of each scenario for subsequent impact assessments. In particular, regional eco-driving effectiveness $E(\cdot)$ is used to assess the overall benefit level as per Equation 6. Each scenario is simulated five times for statistical significance, and results are averaged following the recommended practice (1).
Remark: As we listed in each of the main sections, and more broadly in section 10.13 in the Appendix, we make certain assumptions in our modeling efforts. Hence, the analyses presented here should be interpreted within the framework of these assumptions.
@@ 7.1 Experimental setup
We use Proximal Policy Optimization (PPO) (43) algorithm in training our MT-DRL policies with actor-critic architecture. The actor and critic do not share parameters and are separate neural networks. Centralized training and a decentralized execution paradigm are used for training. Section 10.8 in the Appendix describes the architecture details and all the hyperparameter configurations used in training.
@@ 7.2 Impact of eco-driving
Our eco-driving assessment encompasses diverse perspectives. First, we analyze the overall regional eco-driving effectiveness and the resultant variations in the intersection throughput in the three cities. Second, we analyze which intersection-related factors have a greater influence at different levels of eco-driving adoption. Third, we analyze the level of intersection compatibility needed to achieve the benefits, informing implementation plans. Fourth, we analyze eco-driving behaviors and examine the safety implications of eco-driving, an aspect that has not received adequate attention in previous work (7). Last, we contextualize eco-driving by analyzing its role in joint consideration with travel growth projections and decarbonization technologies such as vehicle electrification and the adoption of hybrid vehicles.
7.2.1 Overall effectiveness
Figure 8: Left Annual average emission reduction vs. eco-driving adoption (by factoring all seasons and peak and off-peak hours). Right Throughput improvement with varying adoption levels. By design, the throughput improvement is at least 0% when evaluating the emission reduction. Also, we ensure these throughput increases do not negatively affect the network-level traffic.
Figures 8 show the annual average effectiveness of eco-driving, assessing the percentage emission reduction and average throughput improvement. On the policy side, throughput gives an alternative argument supporting eco-driving. Eco-driving achieves a notable 11-22% emission reduction, equivalent to the national carbon emissions of Israel and Nigeria, respectively (refer to Appendix 10.2 for the estimation). Notably, a 10% eco-driving adoption can yield 25-50% of the total emission reduction, showcasing a non-linear benefit scaling and greater potential for short-term deployments to gain a majority of total benefits. Importantly, these gains incur no costs to intersection throughput. Additionally, Figure 9 demonstrates consistent emission benefits under different weather conditions across seasons, with minor fluctuations. This emphasizes that eco-driving can yield benefits throughout the year. In the Appendix 10.9, we also show the peak vs off-peak emission benefit percentages.
Figure 9: Impact of weather on eco-driving benefits. For each city and season, we show emission benefits under sunny and rainy conditions. Some curves are offset by ±2% on the x-axis for clarity. Not all weather types occur in every city-season pair.
@@ 7.2.2 Influential factors in eco-driving
Through a Pearson correlation analysis in Figure 11, we underscore key factors influencing emission benefits for informed intersection design and city planning. We consider nine key eco-driving factors that are known to have a significant impact on emission benefits and are directly observable, including temperature, humidity, speed limit, road grade, signal time ratio, vehicle inflow, phase count, lane count, and lane length. The signal ratio denotes the ratio between approach-related phase times and total cycle time. Notably, speed limits, vehicle inflow, and lane lengths significantly impact emission benefits. In particular, when the speed limit and lane length increase, the benefit levels increase. On the other hand, when the inflow increases, the benefit decreases. This explains why the three cities we consider have different benefit levels. As discussed in Appendix Section 10.5, different cities have different intersection feature distributions (refer to Figure 17 in the Appendix). Thus, one plausible explanation is that cities like Atlanta demonstrate higher benefits, given their higher speed limits and longer incoming approaches, compared to relatively low-benefit cities like San Francisco. When the cities are denser (short incoming approaches), there is less opportunity for eco-driving, and when the speed limits are high, there is more room for vehicle speed changes. However, we note that while this is a plausible explanation, the final benefit levels are determined by the combined influence of multiple interacting factors that we consider. We also note that Atlanta, San Francisco, and Los Angeles typically experience relatively mild variations in average temperature and humidity across the four seasons. Since our analysis is based solely on seasonal average temperatures, this limited variability may explain why these factors do not show a significant influence. We also provide a statistical significance test for the analysis performed in this section in Appendix 10.10.
@@
7.2.3 Where to implement eco-driving
@@
In this analysis, we identify which intersections hold the most promise for eco-driving. This is a key question, as enabling eco-driving may require the installation of roadside units at intersections (for V2I communication). Prioritizing intersections that offer the greatest potential benefits can support a more sustainable and strategic deployment plan.
Figure 12 shows Pareto charts for all three cities, revealing that at least 70% of emission benefits can be achieved by implementing eco-driving in just 20% of intersections at each adoption level. This emphasizes the potential for an efficient phased deployment, focusing on a specific subset of intersections. However, the Venn diagram in Figure 13 illustrates the specific 20% of intersections delivering the most significant emission benefits at every adoption level may have little overlap between eco-driving adoption levels. It highlights that intersections effective at lower eco-driving adoption levels may not maintain the same efficacy as eco-driving adoption increases, emphasizing the need for careful deployment planning, such as which intersections to be embedded with roadside units (RSU) for Signal and Phase Timing (SPaT) communication. In broad terms, this would question what infrastructure is considered long-term value for eco-driving, as transportation agencies are often constrained by monetary budgets for infrastructure.
@@ 7.2.4 Eco-driving behaviors
Figure 14: Time-space diagrams of semi-autonomous vehicles (highlighted) and human-driven vehicles (muted) for a selected representative two-lane intersection. The vehicles are driven from 200m relative position ot -200m relative position.
We further analyze the induced road safety metrics of these learned behaviors compared to human-like driving. Figure 15 presents a roadway safety analysis using widely used surrogate safety measures (SSMs) commonly employed in assessing roadway safety (52). We analyze five SSMs, namely: time to collision (TTC), post-encroachment time (PET), maximum speed (MaxS), change in speed (DeltaS), and deceleration rate (DR). Further details on the analysis and used surrogate measures are in Appendix Section 10.11. Results in Figure 15 suggest that the learned behaviors are as safe as human-like driving, resulting from smoother vehicle behaviors. However, given that eco-driving behaviors of controlled vehicles are emerging behaviors that may come as counterintuitive for regular human drivers, it may trigger unforeseen human-driver responses, requiring additional human-factor studies for a better understanding of such human factors at play (22).
@@
Figure 16: Eco-driving benefits in the presence of vehicle electrification and hybrid engines as alternative decarbonization strategies. Figure shows total reductions over 2024-2050 in carbon emissions at various cumulative average eco-driving adoption levels amidst projected electric vehicle and hybrid adoption in each city. Eco-driving yields additional benefits relative to a future with hybridization or electrification alone. For example, with a cumulative average total adoption rate of 20% from 2024 to 2050, eco-driving alone would cut San Francisco emissions in that period by about 7%, and when combined with projected electric vehicle and hybrid adoption, would cut emissions by about 17%. Results are relative to an all internal combustion engine vehicles baseline (i.e., without electrification or hybrid adoption.)
@@ 9 Acknowledgments
This research was partially supported by the Utah Department of Transportation (Project F-ST99(783)), the National Science Foundation (NSF) CAREER award (#2239566), and an Amazon AWS Machine Learning Research Award. The authors extend their appreciation to Michael Sheffield, Christopher Siavrakas, and Kelly Njord at the Utah Department of Transportation for constructive discussions and providing their expert opinions on the work. Authors also extend their appreciation to Steven Barrett, Samitha Samaranayake, Zuduo Zheng, Yashar Farid, Jinhua Zhao, and Felice Frankel for their constructive comments. The authors also thank Sunera Avinash, Anirudh Valiveru, Jiaxin He, and Dajiang Suo for their help in numerous stages of the project. The authors acknowledge MIT SuperCloud and the Lincoln Laboratory Supercomputing Center for providing computational resources supporting the research results in this paper.